# GNN-DT: A Graph Neural Network Enhanced Decision Transformer for Efficient Optimization in Dynamic Environments

### v2

@@ Appendix A Appendix: EV Charging Model
$\omega^{+}_{i,t},\omega^{-}_{i,t}$ Binary operation indicators Indicates if charger $i$ charges ($+$) or discharges (--) at timestep $t$
@@
The first appendix Section provides the complete mixed integer programming (MIP) and MDP formulation of the optimal EV charging problem and presents detailed experimental results for key evaluation metrics.
A.1 Complete EV charging MIP model
The optimal EV charging problem we investigate aims to maximize the CPO’s profits while ensuring that the demands of EV users are fully satisfied. The CPO lacks prior knowledge of when EVs will arrive or the aggregated power constraints. However, it is assumed that when an EV arrives at charging station $i$ initiating charging session $j$, it provides its departure time ($t_{j,i}^{d}$) and the desired battery capacity ($e^{*}_{j,i}$) at departure. Additionally, the battery capacity $e_{i,t}$ for each EV is known while it is connected to the charger. These assumptions are standard in research, as this information can typically be retrieved through evolving charging communication protocols.
The investigated EV charging problem (See notation Table 5) expands over a simulation with $t$ timesteps, $t\in\mathcal{T}$ where a CPO decides the charging and discharging power ($p^{+}_{i,t}$ and $p^{-}_{i,t}$) for every charging station $i\in\mathcal{I}$. Since the chargers can be spread around the city, there are charger groups $w\in\mathcal{W}$, that can have a lower-level aggregated power limits representing connections to local power transformers. Based on these factors, the overall optimization objective is defined as follows:
$\max_{p^{+},\omega^{+},p^{-},\omega^{-}}\sum_{t\in\mathcal{T}}\Bigg{(}$ $-100\cdot\max\{0,p_{t}^{\sum}-p^{*}_{t}\}$ (7)
$+\sum_{i\in\mathcal{I}}\Bigg{(}\Delta t\left(\Pi^{+}_{t}p_{i,t}^{+}\omega^{+}_{i,t}-\Pi^{-}_{t}p_{i,t}^{-}\omega^{-}_{i,t}\right)$ (8)
$\quad-10\cdot\sum_{j\in\mathcal{J}_{i}}\left(\sum_{t=t_{j,i}^{a}}^{t_{j,i}^{d}}\left(p_{i,t}^{+}\omega^{+}_{i,t}-p_{i,t}^{-}\omega^{-}_{i,t}\right)-e^{*}_{j,i}\right)^{2}\Bigg{)}\Bigg{)}$ (9)
@@ A.1 Complete EV charging MIP model
$p^{\sum}_{t}=\sum_{i\in\mathcal{I}}p^{+}_{i,t}\cdot\omega^{+}_{i,t}-p^{-}_{i,t}\cdot\omega^{-}_{i,t}$ $\forall i,\;\forall t$ (10)
@@
$e_{j,i,t}=e_{j,i,t-1}+(p^{+}_{i,t}\cdot\omega^{+}_{i,t}+p^{-}_{i,t}\cdot\omega^{-}_{i,t})\cdot\Delta t$ $\forall j,\;\forall i,\;\forall t$ (13)
@@
The power of a single charger $i$ is modeled using four decision variables, $p^{+}\cdot\omega^{+}$ and $p^{-}\cdot\omega^{-}$, where $\omega^{+}$ and $\omega^{-}$ are binary variables, to differentiate between charging and discharging behaviors and enable charging power to get values in ranges $0\cup[\underline{p}^{+},\overline{p}^{+}]$, and discharging power in $[\underline{p}^{-},\overline{p}^{-}]\cup 0$. Eq. 11 defines the locally aggregated transformer power limits $\overline{p}$ for chargers belonging to groups $\mathcal{W}_{i}$. Eqs.(13) and(14) address EV battery constraints during operation with a minimum and maximum capacity of $\underline{e}$, $\overline{e}$, and energy $e^{a}$ at time of arrival $t^{a}$. Equations (15) and (16) impose charging and discharging power limits for every charger-EV session combination. To prevent simultaneous charging and discharging, the binary variables $\omega^{\text{ch}}$ and $\omega^{\text{dis}}$ are constrained by (17).
A.2 Evaluation Metrics
The following evaluation metrics are used in this study to assess the performance of the proposed algorithms:
User Satisfaction [%]: This metric measures how closely the state of charge (SoC) of an EV at departure matches its target $\text{SoC}^{*}$. For a set of EVs $\mathcal{J}$, user satisfaction is given by:
[…]
This ensures that each EV is charged to its desired level by the end of the charging session.
Energy Charged [kWh]: This represents the total energy supplied to EVs over the entire charging period and is given by:
Energy $\textbf{Energy Charged}=\sum_{t\in\mathcal{T}}\sum_{i\in\mathcal{I}}p^{+}_{i,t}\cdot\omega^{+}_{i,t}\cdot\Delta t.$ (19)
This metric helps quantify the overall energy throughput for the system.
Energy Discharged [kWh]: This measures the amount of energy discharged back to the grid by the EVs. Discharging is typically done when electricity prices are high, and this metric is important for evaluating the system’s ability to contribute to grid stability.
Energy $\textbf{Energy Discharged}=\sum_{t\in\mathcal{T}}\sum_{i\in\mathcal{I}}p^{-}_{i,t}\cdot\omega^{-}_{i,t}\cdot\Delta t.$ (20)
Power Violation [kW]: This metric tracks violations of operational constraints, such as exceeding the aggregated power limits at any given time. A violation occurs when the total power used exceeds the maximum allowed power:
Power $\textbf{Power Violation}=\sum_{t\in\mathcal{T}}\max\{0,p_{t}^{\sum}-p^{*}_{t}\}\Delta t.$ (21)
Minimizing this metric ensures that the system remains within operational limits and avoids overloading the grid.
Cost [€]: This evaluates the financial cost of the charging operations, considering both charging and discharging costs based on electricity prices. The total cost over the simulation period is defined as:
$\textbf{Cost}=\sum_{t\in\mathcal{T}}\sum_{i\in\mathcal{I}}\Delta t(\Pi^{+}_{t}p_{i,t}^{+}-\Pi^{-}_{t}p_{i,t}^{-}),$ (22)
This metric helps assess the cost-effectiveness of the charging strategy.
A.3 The MDP State Space
Figure 5: The EV charging problem graph structure and state. (a) shows the raw physical layout of EVs, chargers and transformer with information and power links; (b) recasts this as a four-type node graph (EV, charger, transformer, CPO) carrying only the observations available to the CPO; and (c) prunes any charger without connected EV resulting in a compact dynamic heterogeneous state graph.
At each time step $t$, the environment state $\boldsymbol{s}_{t}\in\mathcal{S}$ is encoded as a heterogeneous graph $\mathcal{G}_{t}=(\mathcal{N}_{t},\mathcal{E}_{t})$, where $\mathcal{N}_{t}$ is the set of nodes (EVs, chargers, transformer, and CPO) and $\mathcal{E}_{t}$ the set of edges encoding physical and informational connections. Each node $n\in\mathcal{N}_{t}$ carries a feature vector $\boldsymbol{x}^{t}_{n}\in\mathbb{R}^{d}$, encapsulating all observations available to the CPO under V2G protocols. Fig. 5 shows how the EV charging problem is modelled as a graph with each node type having different number of features as detailed below.
An EV node’s feature vector:
@@ A.3 The MDP State Space
encodes its current state of charge $\text{SoC}_{t}\in[0,1]$, the remaining steps until departure $t^{\rm d}-t$, and the identifiers $j$, $i$, and $w$ of the charging point, charger, and transformer group to which it is connected. Exact SOC and departure-time information are assumed available via ISO 15118-style V2G signaling. A charging station (cs) node $i$ is described by:
@@
where $\overline{p}_{w,t}$ denotes its maximum power capacity and $w$ its identifier, ensuring the policy respects feeder-level constraints. The CPO (system) node aggregates temporal and economic context in
@@
Here $d$ and $h$ encode day-of-week and hour to capture demand patterns, $\Pi^{+}_{t}$ is the charging price at $t$, and $p^{\sum}_{t-1}$ is the previous time step’s total power draw, providing feedback for profit-maximizing decisions.
A.4 Complete Experimental Results of EV Charging.

### v3

@@ 3 Problem Formulation
$\omega^{+}_{i,t},\omega^{-}_{i,t}$ Binary operation indicators Indicates if charger $i$ charges (++) or discharges (-) at timestep $t$
@@
3.1 Offline RL
Offline RL aims to learn a policy $\pi_{\theta}(a\!\mid\!s)$ that maximizes the expected discounted return $\mathbb{E}\bigl[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})\bigr]$ without additional interactions with the environment Levine et al. (2020). A Markov Decision Process (MDP) is defined by the tuple $(S,A,P,R,\gamma)$, where $S$ is the state space, $A$ the action space, $P$ the transition function, $R$ the reward function, $\gamma\in(0,1]$ the discount factor Sutton and Barto (2018). In the offline setting, a static dataset $\mathcal{D}=\{(s,a,r)\}$, collected by a (potentially suboptimal) policy, is provided. DTs leverage this dataset by treating RL trajectories as sequences, learning to predict actions that maximize returns based on previously collected experiences. A key component in DTs is the return-to-go (RTG), which for a time step $t$ can be defined as: $G_{t}=\sum_{\tau=t}^{T}\gamma^{\tau-t}\,r_{\tau},$ representing the discounted cumulative reward from $t$ until the terminal time $T$. Offline RL is particularly beneficial when real-time exploration is costly or impractical, while sufficient historical data are available.
3.2 The EV Smart Charging Problem
Working closely with a charge point operator (CPO), it was evident that existing heuristic and mathematical programming charging strategies don’t scale efficiently as EV fleets grow. To address this, we designed the state–action space and objectives around real‐world operational constraints and assumptions provided by the CPO. We consider a set of $\mathcal{I}$ charging stations indexed i, all assumed to be controlled by a CPO over a time window $\mathcal{T}$, divided into $T$ non-overlapping intervals. Since the chargers can be spread around the city, there are charger groups $w\in\mathcal{W}$, that can have a lower-level aggregated power limits representing connections to local power transformers. For a given time window, each charging station $i$ operates a set of $\mathcal{J}$ non-overlapping charging sessions, denoted by $\mathcal{J}_{i}=\{j_{1,i},\cdots,j_{J_{i},i}\}$, where $j_{j,i}$ represents the $j^{th}$ charging event at the $i^{th}$ charging station and $J_{i}=|\mathcal{J}_{i}|$ is the total number of charging sessions seen by charging station i in an episode. A charging session is then represented as $j_{j,i}:\{t^{a}_{j,i},t^{d}_{j,i},\bar{p}_{j,i},e^{*}_{j,i}\},~\forall j,i$, where $t^{a},t^{d},\bar{p}$ and $e^{*}$ represent the arrival time, departure time, maximum charging power, and the desired battery energy level at the departure time. The primary goal is to minimize the total energy cost given by:
Maintaining the desired battery charge at departure is crucial for EV user satisfaction. We model this behavior as:
Eq. (3) defines a sparse reward added at each EV departure based on its departure energy level. Building on the objective functions described by Eqs. (1)-(3), the overall EV charging problem is formulated as a mixed integer programming (MIP) problem, subject to lower-level operational constraints (e.g., EV battery, power levels) as detailed below:
[…]
$\hskip-14.22636pt+\sum_{i\in\mathcal{I}}\Big(\Delta t\big(\Pi_{t}^{+}p_{i,t}^{+}\omega_{i,t}^{+}-\Pi_{t}^{-}p_{i,t}^{-}\omega_{i,t}^{-}\big)$
[…]
@@ 3.2 The EV Smart Charging Problem
$e_{j,i,t}=e_{j,i,t-1}+(p^{+}_{i,t}\cdot\omega^{+}_{i,t}+p^{-}_{i,t}\cdot\omega^{-}_{i,t})\cdot\Delta t$ $\hskip-2.84526pt\forall j,\;\forall i,\;\forall t$ (7)
@@
The multi-objective optimization function in Eq.(4) integrates Eqs. (1)–(3) using experimentally determined coefficients based on practical importance. The power of a single charger $i$ is modeled using four decision variables, $p^{+}\cdot\omega^{+}$ and $p^{-}\cdot\omega^{-}$, where $\omega^{+}$ and $\omega^{-}$ are binary variables, to differentiate between charging and discharging behaviors and enable charging power to get values in ranges $0\cup[\underline{p}^{+},\overline{p}^{+}]$, and discharging power in $\underline{p}^{-},\overline{p}^{-}]\cup 0$. Eq. (5) defines the locally aggregated transformer power limits $\overline{p}$ for chargers belonging to groups $\mathcal{W}_{i}$. Eqs. (6)-(8) address EV battery constraints during operation with a minimum and maximum capacity of $\underline{e}$, $\overline{e}$, and energy $e^{a}$ at time of arrival $t^{a}$. Equations (9) and (10) impose charging and discharging power limits for every charger-EV session combination. To prevent simultaneous charging and discharging, the binary variables $\omega^{\text{ch}}$ and $\omega^{\text{dis}}$ are constrained by (11).
3.3 EV Charging MDP
The optimal EV charging problem can be framed as an MDP: $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},R)$, where $S$ is the state space, $A$ is the action space, $P$ is the transition probability function, and $R$ is the reward function. At any time step $t$, the state $\boldsymbol{s}_{t}\in\mathcal{S}$ is represented by a dynamic graph $\mathcal{G}_{t}=(\mathcal{N}_{t},\mathcal{E}_{t})$, where $\mathcal{N}_{t}$ is the set of nodes and $\mathcal{E}_{t}$ is the set of edges. The graph is dynamic since the number of nodes in the state and action graph can vary in each step, because of EVs’ arrival and departures. Each node $n\in\mathcal{N}_{t}$ has a feature vector $\boldsymbol{x}_{n,t}\in\mathbb{R}^{d}$, capturing node-dependent information such as power limits and prices. This graph structure Orfanoudakis et al. (2025b) efficiently models evolving relationships among EVs, chargers, and the grid infrastructure. In detail, an EV node’s feature vector:
@@ 3.3 EV Charging MDP
encodes its current state of charge $\text{SoC}_{t}\in[0,1]$, the remaining steps until departure $t^{\rm d}-t$, and the identifiers $j$, $i$, and $w$ of the charging point, charger, and transformer group to which it is connected. Exact SoC and departure‐time information is assumed to be available via ISO 15118 V2G signaling. A charging station (cs) node $i$ is described by:
@@
where $\overline{p}_{w,t}$ denotes its maximum power capacity and $w$ its identifier, ensuring the policy respects feeder‐level constraints. The CPO (system) node aggregates temporal and economic context in
@@
Here $d$ and $h$ encode day‐of‐week and hour to capture demand patterns, $\Pi^{+}_{t}$ is the charging price at $t$, and $p^{\sum}_{t-1}$ is the previous time step’s total power draw, providing feedback for profit‐maximizing decisions.
The action space $\mathbf{a}_{t}\in\mathcal{A}$ is represented by a dynamic graph $\mathcal{G}^{\mathbf{a}}_{t}=(\mathcal{N}^{\mathbf{a}}_{t},\mathcal{E}^{\mathbf{a}}_{t})$, where nodes $\mathcal{N}^{\mathbf{a}}_{t}$ correspond to the decision variables of the optimization problem (e.g., EVs). Each node $n\in\mathcal{N}^{\mathbf{a}}_{t}$ represents a single action $a_{i,t}\in\mathbf{a}_{t}$, scaled by the corresponding charger’s maximum power limit. For charging, $a_{i,t}\in[0,1]$, and for discharging, $a_{i,t}\in[-1,0)$. The transition function $\mathcal{P}(\boldsymbol{s}_{t+1}\mid\boldsymbol{s}_{t},\mathbf{a}_{t})$ accounts for uncertainties in EV arrivals, departures, energy demands, and grid fluctuations. Finally, the reward function is the same with the objective described in Eq. (4) and is defined as $R(\boldsymbol{s}_{t},\mathbf{a}_{t})=f_{1}-100f_{2}-10f_{3}$ for a single timestep ($\mathcal{T}=\{t\}$). The reward is guiding the policy to maximize cost savings, respect operational constraints, and meet EV driver requirements. While Eqs.(1) and (2) represent individual EV rewards and aggregated EV penalties, respectively. Eq.3 introduces a sparse reward that activates only when an EV departs, thereby creating complex temporal dependencies. The reward function weights were selected after comprehensive experimentation, ensuring the optimal solution of the MIP remains the same while the RL training is balanced between exploration and penalization.
4 GNN-based Decision Transformer
4.1 Sequence Embeddings
In detail, the State Embedder consists of $L$ consecutive Graph Convolutional Network (GCN) Kipf and Welling (2016) layers, which aggregate information from neighboring nodes as follows:
4.2 Decoding Actions
4.3 Training: Action Masking and Loss Function
The proposed GNN-DT model is trained via supervised learning using an offline trajectory dataset Chen et al. (2021), similarly to offline RL. Specifically, the GPT-2 model is initialized with its default pre-trained weights, which are subsequently fine-tuned end-to-end for the EV charging optimization task. In GNN-DT, the learning of infeasible actions, such as charging an unavailable EV, is avoided through action masking. At each time step $t$, a mask vector $\mathbf{m}_{t}$, which has the same dimension as $\mathbf{a}_{t}$, is generated with zeros marking invalid actions and ones marking valid actions. For example, an action is invalid when the $a_{i,t}\neq 0$ and no EV is connected at charger $i$. The mean squared error between the predicted actions $\widehat{\mathbf{a}}_{t}$ and ground-truth actions $\mathbf{a}_{t}$ from expert or offline trajectories is employed as the loss function. For a window of length $K$ ending at time $t$, training loss is defined as:
5 Experimental Setup
The dataset generation and the evaluation experiments are conducted using the EV2Gym simulator Orfanoudakis et al. (2025a), which leverages real-world data distributions, including EV arrivals, EV specifications, electricity prices, etc. This setup ensures a realistic environment where the state and action spaces accurately reflect real charging stations’ operational complexity. A scenario with 25 chargers is chosen, allowing up to 25 EVs to be connected simultaneously. In this configuration, the action vector has up to 25 variables (one per EV), while the state vector contains around 150 variables describing EV statuses, charger conditions, power transformer constraints, and broader environmental factors. Consequently, the resulting optimization problem is in the moderate-to-large scale range, reflecting the key complexities of real-world EV charging. Each training procedure is repeated 10 times with distinct random seeds to ensure statistically robust findings. All reported rewards represent the average performance over 50 evaluation scenarios, each featuring different configurations (electricity prices, EV behavior, power limits, etc.). Training was carried out on an NVIDIA A10 GPU paired with 11 CPU cores and 80 GB of RAM, using the AdamW optimizer and a LambdaLR scheduler. Baseline RL agents converged in 2–5 hours, while the proposed GNN-DT required up to 10 hours of training. Default hyperparameters were used for all baseline RL methods; Table 2 lists the full set of hyperparameters employed to train the DTs.
Table 2: Algorithm hyperparameters for small- and large-scale settings.
Hyperparameter Small scale Large scale
Batch size 128 64
Learning rate $10^{-4}$ $10^{-4}$
Weight decay $10^{-4}$ $10^{-4}$
Steps per iteration 1000 3000
Decoder layers 3 3
Attention heads 4 4
Embedding dimension 128 256
GNN embedder feat. dim. 16 16
GNN hidden dimension 32 64
GCN layers 3 3
Epochs 250 400
CPU memory (GB) 8 40
Time limit (h) 10 46
5.1 Dataset Generation
Offline RL algorithms, including DTs, can learn policies from trajectories without needing online interaction with the environment. Consequently, the quality of the gathered training trajectories has a substantial impact on the learning process. In this work, three distinct strategies were used to generate trajectories:
  • Random Actions: Uniformly sampled actions in the range [1,1][-1,1] were applied to the simulator.
  • Business-as-Usual (BaU): A Round Robin charging policy commonly employed by CPOs, which sequentially allocates charging power among EVs to balance fairness and efficiency.
  • Optimal Policies: Optimal solutions derived from solving offline the mathematical problem described in Section 3.2 for randomly generated scenarios.
Each trajectory consists of 300 state-action-reward-action mask tuples, with each timestep representing a 15-minute interval, resulting in a total of three simulated days. This combination of random, typical, and expert data provides a comprehensive basis for evaluating how GNN-DT learns from diverse offline trajectories.
6 Experiments
6.1 Training Performance
Fig. 2 compares the proposed GNN-DT against multiple baselines, including the classic DT Chen et al. (2021) and Q-DT Hu et al. (2024) , which both rely on flattened state representations due to their inability to directly process graph-structured data. In these baseline methods, empty chargers and unavailable actions are replaced by zeros, so the action vector is always the same size. Several well-known online RL algorithms from the Stable-Baselines-3 Raffin et al. (2021) framework are evaluated, such as SAC, DDPG, Twin Delayed DDPG (TD3), Trust Region Policy Optimization (TRPO), Proximal Policy Optimization (PPO), and Truncated Quantile Critics (TQC). Also, offline RL algorithms from D3RLPY Seno and Imai (2022), namely Implicit Q-Learning (IQL), Conservative Q-Learning (CQL), and Behavioral Cloning (BC), are also included. The offline RL algorithms (IQL, CQL, BC, DT, Q-DT, and GNN-DT) are trained on three datasets (Optimal, Random, and BaU), each comprising 10.000 trajectories. A red dotted line marks the optimal reward, which represents the experimental maximum achievable reward obtained by solving the deterministic MIP knowing the future (Oracle) defined in Eq. 4. This oracle reward serves as an upper bound and helps contextualize the relative performance of each method. It is important to note that the EV smart charging problem requires real-time (1-5 min intervals) optimization solutions for large-scale, highly stochastic scenarios, where metaheuristic algorithms, stochastic optimization, and model predictive control methods fail due to computational constraints. By contrast, once trained (over 2–24 hours), RL agents can deliver real-time charging schedules on an ordinary computer in milliseconds.
Table 3: Comparison of maximum episode rewards ($\times 10^{5}$) for baselines and GNN-DT across datasets and context lengths ($K$). Bold indicates the highest value within each dataset and $K$ category.
Dataset Avg. Training Dataset Reward K=2 K=10
DT Q-DT GNN-DT DT Q-DT GNN-DT
Random $100$ 2.37-2.37 ±$0.39$ 1.91-1.91 1.97-1.97 $\mathbf{-0.82}$ 2.12-2.12 2.09-2.09 1.16-1.16
Random $1000$ 1.93-1.93 2.04-2.04 0.86-0.86 2.11-2.11 2.01-2.01 1.18-1.18
Random $10000$ 1.76-1.76 2.04-2.04 1.25-1.25 1.81-1.81 1.98-1.98 $\mathbf{-0.98}$
BaU $100$ 0.67-0.67 ±$0.07$ 0.79-0.79 0.74-0.74 $\mathbf{-0.59}$ 0.79-0.79 0.72-0.72 0.56-0.56
BaU $1000$ 0.71-0.71 0.66-0.66 0.65-0.65 0.64-0.64 0.71-0.71 0.57-0.57
BaU $10000$ 0.69-0.69 0.66-0.66 0.66-0.66 $\mathbf{-0.44}$ 0.74-0.74 0.53-0.53
Optimal $100$ 0.01-0.01 ±$0.01$ 0.67-0.67 0.91-0.91 0.15-0.15 1.12-1.12 0.90-0.90 0.14-0.14
Optimal $1000$ 0.63-0.63 0.67-0.67 0.10-0.10 0.87-0.87 0.86-0.86 0.09-0.09
Optimal $10000$ 0.63-0.63 0.80-0.80 $\mathbf{-0.04}$ 0.72-0.72 0.90-0.90 $\mathbf{-0.07}$
In Figs. 2.a-c, the DT-based approaches use a context length $K=10$. As expected, the Optimal dataset provides the highest-quality information, enabling GNN-DT to converge rapidly toward near-oracle performance, while classic DT, Q-DT, and the other offline RL algorithms lag far behind, showcasing GNN-DT’s improved sample efficiency. With the Random dataset, the limited quality of data leads all methods to plateau at lower reward values, although GNN-DT still surpasses the other baselines. An intriguing behavior is observed with the BaU dataset, where classic DT, BC, and IQL converge at rewards exceeding those of GNN-DT. In contrast, the online RL algorithms displayed in Fig. 2.d struggle to achieve comparable improvements, suggesting that pure online exploration is insufficient for solving this complex EV charging optimization problem with sparse rewards. In the rest of this section, we omit the online and offline RL baselines, as their performance is substantially inferior to that of DT, Q-DT, and GNN-DT.
6.2 Dataset Impact
In Table 6.1, the maximum episode reward is compared for small, medium, and large datasets (100, 1.000, and 10.000 trajectories), under two different context lengths ($K=2$ and $K=10$). The left side of Table 6.1 reports the dataset type, the number of trajectories, and the average reward in each dataset. All baselines achieve performance above the Random dataset’s average reward. However, only GNN-DT consistently approaches the Optimal dataset’s performance, reaching as close as $0.04\times 10^{5}$ compared to the $0.01\times 10^{5}$ optimal reward. This advantage becomes especially evident at the largest dataset size (10.000 trajectories), highlighting the benefits of the graph-based embedding layer. Overall, GNN-DT outperforms the baselines across all datasets and both context lengths, with the single exception of the BaU dataset at $K=10$. Interestingly, a larger context window does not always translate into higher rewards, potentially due to the problem setting. Similarly, the dataset size appears to have minimal impact on Q-DT, whereas DT and GNN-DT generally improve with more trajectories. These findings underscore that both the quality and quantity of offline data, coupled with the GNN-DT architecture, are key to achieving superior performance.
6.3 Enhancing Training Datasets
The previous section highlighted that the quality of trajectories in the training dataset is the most influential factor for achieving high performance. In this section, we explore whether creating new datasets by mixing existing ones can further improve performance. The Optimal and Random datasets are combined in different proportions, as summarized in Table 4. A noteworthy result is that supplementing the Optimal dataset with “less useful” (Random) trajectories consistently boosts performance. In particular, GNN-DT with $K=10$, trained on a mix of 250 Optimal and 750 Random trajectories, achieves near-oracle results, deviating by only $0.001\times 10^{5}$ from the optimal reward. A similar trend emerges when blending BaU and Random datasets shown in Table 5. While the BaU dataset alone performs worse than the Optimal dataset, mixing it with Random data still yields improvements, with the 75% BaU and 25% Random combination showing the best results. Overall, these findings indicate that carefully integrating high- and lower-quality data can enhance policy learning beyond what purely Optimal or purely Random datasets can provide.
Table 4: Maximum reward of GNN-DT trained on merged Optimal and Random datasets for $K=2$ and $K=10$. Performance improves despite lower average training rewards, highlighting the importance of dataset diversity. Highest rewards per $K$ are highlighted with bold.
Dataset Total Traj. Avg. Dataset Reward GNN-DT Reward ($\times 10^{5}$)
K=2 K=10
Random (Rnd.) 100% $1000$ 2.37-2.37 ±$0.39$ 0.863-0.863 1.187-1.187
Opt. 25% + Rnd. 75% $1000$ 1.78-1.78 ±$1.07$ 0.045-0.045 $\mathbf{-0.020}$
Opt. 50% + Rnd. 50% $1000$ 1.18-1.18 ±$1.19$ $\mathbf{-0.021}$ 0.040-0.040
Opt. 75% + Rnd. 25% $1000$ 0.60-0.60~±$1.03$ 0.073-0.073 0.057-0.057
Optimal (Opt.) 100% $1000$ 0.01-0.01 ±$0.01$ 0.108-0.108 0.099-0.099
Dataset Total Traj. Avg. Dataset Reward GNN-DT Reward ($\times 10^{5}$)
K=2 k=10
Random (Rnd.) 100% $1000$ 2.37-2.37 ±$0.39$ 0.863-0.863 1.187-1.187
BaU 25% + Rnd. 75% $1000$ 1.93-1.93 ±$0.80$ 0.578-0.578 0.461-0.461
BaU 50% + Rnd. 50% $1000$ 1.51-1.51 ±$0.87$ 0.665-0.665 $\mathbf{-0.447}$
BaU 75% + Rnd. 25% $1000$ 1.09-1.09 ±$0.76$ $\mathbf{-0.421}$ 0.471-0.471
BaU 100% $1000$ 0.01-0.01 $\textpm 0.01$ 0.654-0.654 0.572-0.572
Figure 3: GNN-DT performance for larger context lengths (K).
DT State GNN Action GNN Res. Con. Action Mask Optimal ($\times 10^{5}$) Mixed ($\times 10^{5}$)
$0.69\pm 0.03$ $0.95\pm 0.39$
GCN $1\pm 0.02 0.77±0$$7\pm 0.17$
GCN $0.18\pm 0.03$ $0.16\pm 0.07$
GCN GCN $0.11\pm 0.03$ $0.12\pm 0.04$
GAT GAT $0.14\pm 0.07$ $0.15\pm 0.06$
GCN GCN $\mathbf{-0.09\pm 0.02}$ $\mathbf{-0.10\pm 0.04}$
6.4 Impact of larger context lengths (K)
6.5 Component Ablation Study
To better understand the contribution of each architectural component, we conduct an ablation study that systematically removes or replaces elements of our model and then trains the model on the Optimal and Mixed (Opt.25% +Rand.75%). The results in Table 6 reveal that neither a plain DT nor a DT augmented solely with a state-GNN submodule achieves competitive performance. Notably, adding the residual connection atop the state-GNN leads to a significant improvement, from $0.77\times 10^{5}$ to $0.16\times 10^{5}$ on the Mixed dataset, demonstrating its importance for effective credit assignment over dynamic inputs. Removing action masking or replacing the GCN module with a Graph Attention Network (GAT) Veličković et al. (2018) similarly degrades performance, indicating that each component provides distinct and complementary benefits. Ultimately, only the full GNN-DT architecture achieves strong performance across both the Mixed and Optimal datasets.
6.6 Average Results of EV Charging.
Table 7 shows a comparison of key EV charging metrics for the 25-station problem after 100 evaluations, including heuristic algorithms, Charge As Fast as Possible (CAFAP) and BaU, and DT variants with the optimal solution, which assumes future knowledge.
The performance of the proposed algorithms was assessed using several evaluation metrics. For example, user satisfaction [%] captures the extent to which the state of charge at departure ($e_{j,t^{d}}$) of each electric vehicle $j\in\mathcal{J}$ meets its target $e_{j}^{*}$, thus defined as:
$\text{User Satisfaction [\%]}=\frac{1}{|\mathcal{J}|}\sum_{j\in\mathcal{J}}\left(\frac{e_{j,t^{d}}}{e_{j}^{*}}\right)\cdot 100\%.$ (18)
Energy charged [kWh] was measured as the total amount of energy delivered to the vehicles during the charging sessions, while energy discharged [kWh] was quantified as the energy returned from vehicles to the grid. Power violations [kW] were tracked to identify instances in which operational limits were exceeded, ensuring system feasibility. Finally, the overall charging cost [€] was evaluated by accounting for the time-varying electricity prices during charging and discharging periods, thus reflecting the economic performance of the strategy.